# Are Neuromorphic Architectures Inherently Privacy-preserving? An Exploratory Study

### v1

@@ 6.1. MIA Assessment
In this section, we evaluate the resilience of both ANN and SNN models to MIAs across multiple architectures, including the baseline, ResNet18, and VGG16 models for each dataset. The evaluation metric used is the Area Under the Curve (AUC) of the Receiver Operating Characteristic (ROC) curve, where lower AUC values signify greater resistance to MIAs.
Our findings consistently demonstrate that SNNs achieve lower AUC values than their ANN counterparts across all datasets, highlighting their stronger inherent resilience against privacy attacks. For instance, on the MNIST dataset (Figure 4a), SNN models yield an AUC as low as 0.51, in contrast to ANNs, which range between 0.59 and 0.61. This pattern persists in more complex datasets like CIFAR-10 (Figure 4c) and CIFAR-100 (Figure 4d), where SNNs achieve respective AUCs of 0.53 and 0.58 across model variations, compared to ANNs which reach AUCs up to 0.82 and 0.88, respectively.
As shown in Table 2, across all architectures, SNNs not only maintain a robust classification accuracy but also exhibit significantly lower vulnerability to MIAs, suggesting an advantage in privacy-preserving properties without compromising performance. The contrast is particularly pronounced in high-dimensional data like CIFAR-100, where ANNs display considerable susceptibility to MIAs (AUC values nearing 0.88), whereas SNNs consistently maintain lower AUCs across all model architectures.
Is SNN resilience driven by ANN overfitting?
To assess whether the improved resilience of SNNs is simply a result of ANN overfitting, we examined both ANN and SNN models trained on the baseline architecture across multiple datasets, including F-MNIST, Breast Cancer, and CIFAR-10. Overfitting is known to increase a model’s vulnerability to MIA by reducing the generalizability of the model. Overfitted models tend to memorize training data, which increases the likelihood that an adversary can successfully infer whether a particular data point was part of the training set, thus leading to a higher MIA AUC.
@@
The pronounced overfitting in CIFAR-10 (Figure 5c) can be attributed to the dataset’s complexity and the relatively shallow two-convolution-layer architecture, which, over extended training epochs, struggles to generalize effectively and tends to memorize the training data. Despite these similar overfitting trends, SNNs consistently exhibit lower MIA AUCs than their ANN counterparts. This suggests that the observed resilience of SNNs against MIAs is not due to overfitting in ANNs, but rather reflects inherent differences in how SNNs process and encode information, making them less susceptible to privacy leakage.
@@ 6.2. SNN Exploration Space Assessment:
In the SNN domain, we analyze the results by examining different learning algorithms, the frameworks utilized for these algorithms, and the specific parameters within each framework.
@@ 6.2.1. Surrogate Gradient Algorithm
Impact of Encoding Schemes in snnTorch Framework:
Figure 6 illustrates the influence of the three encoding schemes available in snnTorch on the vulnerability of SNNs to MIA. For the MNIST (Figure 6a) and F-MNIST (Figure 6b) datasets, Rate and Delta encoding schemes yield similar AUC values, approximately 0.51 and 0.53, respectively, suggesting a marginal effect on MIA susceptibility. By contrast, the CIFAR-10 dataset (Figure 6c) shows higher AUC values across all encoding types, with values reaching up to 0.58, indicating a slightly elevated vulnerability compared to other datasets. The Iris dataset (Figure 6d) displays more pronounced variation, with Delta encoding achieving a slightly higher AUC of 0.56, suggesting a subtle impact of encoding schemes in smaller datasets. Overall, while no consistent pattern emerges regarding AUC across different encoders, the results consistently fall within a lower AUC range, underscoring the inherent resilience of SNNs towards MIAs across datasets and encoding methods.
@@
In our experimental setup using the LAVA framework, we employed a three-layer feed-forward neural network architecture featuring Leaky Integrate-and-Fire (LIF) neurons. This network was configured with pre-trained synaptic weights.
As Figure 8 demonstrates, the attack on SNN within the LAVA framework results in an AUC of 0.52, which is about 14.75% lower than its ANN (PyTorch) counterpart.
@@
Figure 9. Model Performance over increasing Privacy Budget ($\varepsilon$) with DPSGD across MNIST, F-MNIST, CIFAR-10, Iris, and Breast Cancer datasets.
@@ 6.2.2. Evolutionary Algorithm:
Figure 7a and Figure 7b respectively depict the results from evaluating Evolutionary Optimization parameters within the TennLab Framework on the Breast Cancer and datasets. For Iris, AUCs range from 0.50 to 0.53 across varying population sizes ( 7b(i)), and remain close to 0.50 with different crossover ( 7b(iii)) and mutation rates ( 7b(ii)), indicating consistent protection. Similarly, the Breast Cancer dataset maintains AUCs near 0.50, with minimal impact from changes in population size ( 7a(i)), crossover rate ( 7a(iii)), or processor type ( 7a(iv)) such as RISP, Caspian, or Ravens.
Compared to surrogate gradient learning (Table 2), where Iris and Breast Cancer datasets scored AUCs of 0.57 and 0.55 respectively, evolutionary algorithms consistently achieve lower AUCs, suggesting an enhanced resilience to MIAs. This consistent resilience is particularly notable, as it indicates that evolutionary algorithms may provide a more robust defense against privacy attacks. Given that evolutionary optimization supports practical hardware implementation within the TennLab framework, achieving a notable privacy gain through this approach could offer substantial benefits for secure neuromorphic applications.
@@ 6.3. Privacy-Utility Trade off Assessment
MNIST 0.22 - 2.00 7.89% 6.65%
Fashion-MNIST 0.22 - 2.00 19.55% 12.58%
CIFAR-10 0.22 - 2.00 34.43% 27.87%
Breast Cancer 0.22 - 2.00 6.23% 1.93%
Iris 0.22 - 2.00 29.63% 19.04%
To evaluate the privacy-utility trade-off, we applied Differentially Private Stochastic Gradient Descent (DPSGD) with a fixed privacy budget (epsilon, $\epsilon$) of 0.22-2 across both ANN and SNN models, establishing a consistent basis for comparison. Our results, presented in Figure 9 and Table 3, reveal that SNNs generally experience a smaller drop in classification accuracy compared to ANNs under equivalent privacy constraints. For instance, in the MNIST dataset (Figure 9a), SNN models show an average accuracy drop of 6.65%, while ANNs experience a slightly more pronounced decline of 7.89%. This trend is similarly observed in the F-MNIST dataset (Figure 9b), where SNN accuracy decreases by 12.58% compared to a 19.55% reduction in ANNs. Notably, in CIFAR-10 (Figure 9c), the accuracy drop is 27.87% for SNNs and 34.43% for ANNs, demonstrating that SNNs maintain greater utility even as data complexity increases.
The Breast Cancer dataset (Figure 9e) further illustrates this pattern, with SNNs showing only a 1.93% drop in accuracy compared to a 6.23% decrease in ANNs, reflecting the robustness of SNNs in scenarios with stringent privacy requirements. Similarly, in the Iris dataset (Figure 9d), SNNs exhibit a 19.04% accuracy drop, whereas ANNs experience a 29.63% reduction. These results underscore a consistent advantage for SNNs in balancing privacy and utility, as they retain higher performance levels under equivalent privacy budgets.
@@ 7. Conclusion
Our experimental framework encompasses various learning algorithms and programming frameworks, providing a thorough analysis of the factors influencing the privacy preserving properties of SNNs. Our key findings demonstrate that SNNs have a greater ability to withstand MIAs compared to ANNs. Specifically, evolutionary learning algorithms have proven to be more effective than gradient-based methods in strengthening the privacy resilience of SNNs. Additionally, our analysis of the privacy-utility trade off reveals that SNNs show less performance degradation than ANNs when we apply the DPSGD algorithm.
These findings suggest promising directions for privacy-sensitive applications, particularly in scenarios where efficient hardware implementation is required. However, it should be noted that these advantages in privacy preservation and hardware efficiency are specific to certain applications and do not indicate SNNs’ superiority across all domains.
As we move forward, our research endeavors will concentrate on the deployment of our models on Intel’s Loihi neuromorphic chip, enabling the execution of these privacy attacks in a hardware setting. This will provide valuable insights into the practical implementation and real-world performance of SNNs. Concurrently, we intend to expand the scope of our evaluations by incorporating more advanced and sophisticated attack methodologies, subjecting SNNs to increasingly potent threats to further assess their resilience and identify potential vulnerabilities. Moreover, we are actively investigating the integration of differential privacy techniques through evolutionary algorithms within the TennLab framework, aiming to enhance the data privacy protection capabilities of SNNs and explore novel approaches to mitigate privacy risks in neuromorphic computing systems. These ongoing efforts will contribute to a comprehensive understanding of the privacy preserving properties of SNNs and guide the development of robust and secure neuromorphic architectures for privacy-sensitive applications.

### v2

@@ 5.1. MIA Assessment
This section presents a comprehensive evaluation of ANN and SNN architectural resilience against MIAs across baseline, ResNet18, and VGG16 configurations. The privacy vulnerability is meausured using ROC-AUC metrics, where lower values indicate enhanced privacy preservation. Each experiment was repeated three times, with Table 2 reporting the mean and standard deviation. The low standard deviations across all metrics indicate the statistical stability of the results.
The experimental results, illustrated in Figure 4 and Table 2, demonstrate consistent privacy advantages of SNNs while maintaining competitive accuracy. For MNIST, while both architectures achieve comparable test accuracy ($\simeq 99\%$), SNNs exhibit lower AUC values (0.51-0.52) compared to ANNs (0.59-0.61). This privacy advantage becomes more pronounced in F-MNIST, where SNNs maintain AUC values between 0.53-0.55 while achieving 90-92% test accuracy, compared to ANNs’ higher vulnerability (AUC 0.62-0.67) at similar accuracy levels.
The privacy advantage becomes more pronounced in complex datasets. For CIFAR-10, the baseline SNN maintains an AUC of 0.59 compared to ANN’s 0.82, while ResNet18 and VGG16 SNN variants achieve AUCs of 0.55 and 0.53, respectively, substantially lower than their ANN counterparts (0.63 and 0.67). This trend persists in CIFAR-100, where the baseline SNN demonstrates an AUC of 0.58 versus ANN’s 0.88, with ResNet18 and VGG16 SNN implementations maintaining AUCs of 0.66 and 0.58 compared to ANNs’ 0.77 and 0.70. The ROC curves in Figure 4 visually demonstrate this enhanced privacy preservation across all architectural configurations. Notably, this privacy advantage does not significantly compromise model utility, as evidenced by the competitive accuracy metrics maintained across implementations.
Key Finding 1: SNNs demonstrate enhanced privacy preservation across all architectural configurations compared to ANNs, while maintaining competitive accuracy.
5.2. Is SNN resilience driven by ANN overfitting?
To assess whether the improved resilience of SNNs is simply a result of ANN overfitting, the relationship between architectural privacy advantages and model generalization is examined across multiple datasets, including F-MNIST, Breast Cancer, and CIFAR-10. Overfitting is known to increase a model’s vulnerability to MIA by reducing its generalizability. Overfitted models tend to memorize training data, which increases the likelihood that an adversary can successfully infer whether a particular data point was part of the training set, thus leading to a higher MIA AUC.
@@ 5.2. Is SNN resilience driven by ANN overfitting?
The consistency in overfitting patterns across both architectures demonstrated through comparable generalization gaps indicates that ANN’s increased vulnerability to MIAs cannot be attributed to differential overfitting behavior. Despite similar levels of overfitting, SNNs consistently exhibit lower MIA AUC values compared to ANNs. This finding establishes that the enhanced privacy preservation in SNNs originates from inherent architectural characteristics rather than advantages in generalization capability, as both architectures demonstrate comparable overfitting tendencies while exhibiting distinctly different privacy vulnerabilities.
@@ 5.3. SNN Exploration Space Assessment:
This section examines SNN privacy characteristics across different learning algorithms, frameworks, and their associated parameters.
@@ 5.3.1. Surrogate Gradient Algorithm
Impact of Encoding Schemes in snnTorch Framework: Figure 6 illustrates the differential impact of snnTorch’s three encoding schemes on SNN vulnerability to membership inference attacks. For lower complexity datasets such as MNIST (Figure 6a) and F-MNIST (Figure 6b), Rate and Delta encoding mechanisms demonstrate comparable privacy preservation characteristics, with AUC values converging around 0.51 and 0.53 respectively. The CIFAR-10 dataset (Figure 6c) exhibits a broader vulnerability profile across all encoding implementations, with AUC values extending to 0.58, indicating slightly elevated susceptibility compared to simpler datasets. In the Iris dataset (Figure 6d), the impact of encoding choice becomes more pronounced, with Delta encoding demonstrating marginally higher vulnerability (AUC = 0.56), suggesting increased encoding sensitivity in lower-dimensional datasets. Despite these variations across datasets and encoding methods, the consistently lower range of AUC values underscores the inherent privacy preserving characteristics of SNN architectures.
@@
The LAVA framework implementation employs a sophisticated three-layer feed-forward neural network architecture incorporating Leaky Integrate-and-Fire (LIF) neurons with pre-trained synaptic weights. The experimental results, illustrated in Figure 8, show substantial privacy advantages in the LAVA-based SNN architecture, achieving an AUC of 0.52. This represents a significant 14.75% reduction in vulnerability compared to the conventional PyTorch-based ANN implementation, highlighting the framework’s efficacy in preserving privacy while maintaining neuromorphic computing capabilities.
@@
[…]
Figure 9. Model Performance over increasing Privacy Budget ($\varepsilon$) with DPSGD across (a) MNIST, (b) F-MNIST, (c) CIFAR-10, (d) Iris, and (e) Breast Cancer datasets.
@@ 5.3.2. Evolutionary Algorithm:
Figures 7a and 7b present comprehensive evaluations of Evolutionary Optimization parameters within the TennLab Framework across the Breast Cancer and Iris datasets respectively. The Iris dataset demonstrates robust privacy preservation across multiple parametric variations: population size modifications (Figure 7b(i)) yield AUCs ranging from 0.50 to 0.53, while variations in crossover rates (Figure 7b(iii)) and mutation rates (Figure 7b(ii)) consistently maintain AUCs approaching 0.50, indicating stable privacy protection. The Breast Cancer dataset exhibits similarly robust characteristics, maintaining AUC values near 0.50 across diverse parameters including population size variations (Figure 7a(i)), crossover rate adjustments (Figure 7a(iii)), and different processor implementations (Figure 7a(iv)) such as RISP, Caspian, and Ravens. When compared to surrogate gradient learning implementations (Table 2), where Iris and Breast Cancer datasets demonstrated AUCs of 0.57 and 0.55 respectively, evolutionary algorithms consistently achieve superior privacy preservation with significantly lower AUC values. This enhanced resilience to membership inference attacks is particularly noteworthy given the practical hardware implementation capabilities of evolutionary algorithms within the TennLab framework. The combination of robust privacy preservation and hardware practicality suggests that evolutionary optimization approaches may offer compelling advantages for secure neuromorphic computing applications.
Key Finding 2: Different SNN algorithms (surrogate gradient methods, STDP, and evolutionary optimization) demonstrate consistent privacy preservation across architectures and parameters, with evolutionary algorithms showing particular promise in combining privacy resilience with hardware practicality.
@@ 5.4. Privacy-Utility Trade off Assessment
MNIST 0.22 - 2.00 7.89(±0.14)6.65(±0.08)%
Fashion-MNIST 0.22 - 2.00 19.55(±0.84)12.58(±0.78)%
CIFAR-10 0.22 - 2.00 34.43(±0.20)27.87(±0.11)%
Breast Cancer 0.22 - 2.00 6.23(±0.33)1.93(±0.47)%
Iris 0.22 - 2.00 29.63(±0.81)19.04(±0.72)%
The privacy-utility evaluation implements DPSGD across both architectures with a standardized privacy budget ($\epsilon$) range of 0.22-2. The comparative analysis, presented in Figure 9 and Table 3, quantifies accuracy degradation as the differential between baseline accuracy (pre-DPSGD) and DPSGD implementation accuracy throughout training epochs. The results demonstrate enhanced utility preservation in SNN implementations across all datasets under equivalent privacy constraints. On the MNIST dataset (Figure 9a), SNNs show an average accuracy reduction of 6.65% compared to ANNs’ 7.89%. This pattern extends to F-MNIST (Figure 9b), where SNNs demonstrate a 12.58% accuracy decrease versus ANNs’ 19.55%. CIFAR-10 (Figure 9c) maintains this trend with SNNs showing 27.87% reduction compared to ANNs’ 34.43%, indicating sustained utility preservation even with increased data complexity. The enhanced utility preservation extends to tabular datasets, with Breast Cancer (Figure 9e) showing minimal SNN accuracy degradation of 1.93% compared to ANNs’ 6.23%. Similarly, the Iris dataset (Figure 9d) demonstrates SNN accuracy reduction of 19.04% versus ANNs’ 29.63%. These results highlight SNNs’ consistent ability to maintain utility under privacy constraints across diverse data modalities.
Key Finding 3: Under equivalent differential privacy budget, SNNs consistently demonstrate lower accuracy degradation than ANNs across all datasets, indicating better utility preservation while maintaining privacy guarantees.
6. Related Work
As machine learning systems handle increasingly sensitive data, the potential for privacy violations becomes increasingly significant. Li et al. (Liu et al., 2021a) categorize these privacy challenges into two primary areas: privacy attacks and privacy preserving techniques. Privacy attacks have emerged as a significant concern in ML due to the growing realization that models can inadvertently leak sensitive data. These attacks can broadly be classified into different types, such as model inversion attacks, model extraction attacks, MIA. Model inversion attacks (Fredrikson et al., 2015) reconstruct input data from outputs, while extraction attacks  (Juuti et al., 2019) replicate model’s functionality without direct access to its architecture or parameters. Among these, MIAs are notable for inferring whether a specific data point was used in training. According to the survey conducted by Hu et al. (Hu et al., 2022), MIAs were first proposed in the context of genomic data by Homer et al. (Homer et al., 2008) where an attacker could identify an individual’s genome in a dataset based on summary statistics. Later, Shokri et al.(Shokri et al., 2017) introduced the first systematic MIA framework, showing how adversaries could use shadow models to infer training data membership. Salem et al.  (Salem et al., 2018) reduced the complexity by demonstrating that a single shadow model can perform well compared to using multiple models, and they introduced metric based attacks that rely on confidence scores and entropy without the need for identical data distribution between shadow and target models. Nasr et al. (Nasr et al., 2019) further expanded MIA into white box settings, demonstrating that attackers with access to internal model parameters can perform even more effective MIAs. Melis et al. (Melis et al., 2019) extended MIA to federated learning, highlighting vulnerabilities in distributed learning settings, where multiple parties collaboratively train a model. Song and Mittal (Song and Mittal, 2021) highlighted the increased privacy risks in generative models such as GANs, where membership inference attacks could be carried out on synthetic data generators. Recent work by Ilyas et al. introduced LiRA (Likelihood Ratio Attack)  (Carlini et al., 2022), a method that further improves the accuracy of MIAs by leveraging confidence scores more effectively to distinguish training from non-training data points. In 2024, Zarifzadeh et al. (Zarifzadeh et al., 2024) introduced RMIA, a high-power membership inference attack that outperforms prior methods like LiRA and Attack-R, demonstrating superior robustness, particularly at low false positive rates (FPRs), using likelihood ratio tests.
To counteract these privacy attacks, several privacy preserving techniques have been developed.These techniques range from cryptographic approaches like homomorphic encryption(HE) (Yi et al., 2014) and secure multi-party computation(SMPC) (Zhao et al., 2019) to learning based defenses such as model obfuscation (Guo et al., 2021) and knowledge distillation (Gou et al., 2021). However, these methods often suffer from computational inefficiencies, particularly in large scale systems. Another approach is Federated Learning(FL) (Li et al., 2020), which enables collaborative model training without sharing raw data, but remains vulnerable to attacks like MIA. Among these, Differential Privacy(DP) (Dwork, 2006) has gained prominence due to its strong theoretical guarantees and practical applicability in machine learning settings. It provides a systematic framework for protecting individual data points by introducing noise during computations. In ML, this concept has been adapted through various algorithms, with DPSG) (Song et al., 2013) being the most prominent which applies DP principles by adding noise to the gradient updates during training, offering a practical way to maintain privacy while training large models without significantly compromising accuracy.
While much of the data privacy research has centered around ANNs, expanding these investigations to SNNs is necessary. SNNs not only offer performance levels comparable to ANNs but also exhibit superior energy efficiency and hardware integration capabilities, positioning them as promising candidates for exploring inherent privacy features. Although privacy attacks on neuromorphic architectures remain underexplored, existing studies have yet to confirm SNNs’ potential resistance to such threats. However, significant strides have been made in privacy preserving techniques within the neuromorphic domain. For instance, recent efforts by Han et al.  (Han et al., 2023) focus on developing privacy preserving methods for SNNs, particularly utilizing FL and DP to address both computational efficiency and privacy challenges. Li et al.(Li et al., 2023) introduced a framework that combines Fully Homomorphic Encryption (FHE) with SNNs, enabling encrypted inference while preserving SNNs’ energy efficiency and computational advantages. Similarly, Nikfam et al.(Nikfam et al., 2023) developed an HE framework tailored for SNNs, offering enhanced accuracy over Deep Neural Networks(DNNs) under encryption, while balancing computational efficiency. Additionally, Safronov et al.(Kim et al., 2022) proposed PrivateSNN, a privacy preserving framework for SNNs that employs differential privacy to mitigate membership inference attacks, maintaining the energy efficient nature of SNNs.
@@ 7. Conclusion
The increasing deployment of machine learning systems in privacy sensitive domains has heightened the need for architectures that inherently protect data privacy while maintaining computational efficiency. This investigation addresses these requirements through a thorough examination of privacy characteristics in SNNs, evaluating their resilience against privacy attacks compared with traditional neural architectures. The discrete, event driven nature of spike based processing and temporal dynamics in SNNs may inherently limit information leakage compared to the continuous activations in ANNs, providing natural defense mechanisms against privacy attacks.
The experimental analysis establishes enhanced privacy preservation in SNN architectures, with attack AUC values significantly lower than traditional ANNs across all evaluated datasets (CIFAR-10: 0.59 vs 0.82; CIFAR-100: 0.58 vs 0.88). The privacy gain is particularly pronounced when employing evolutionary learning algorithms, which demonstrate superior resilience compared to gradient-based methods. Additionally, SNNs exhibit improved utility preservation under differential privacy constraints, maintaining higher accuracy levels compared to ANNs when implementing DPSGD across diverse datasets.
While these findings highlight SNNs’ potential for privacy sensitive applications, particularly in resource constrained environments, they are focused on privacy preservation applications. Despite their privacy advantages, SNNs face challenges including complex training processes, potential scalability limitations, and and reliance on specialized hardware, which is necessary for optimal performance. However, within the scope of privacy preservation, their unique computational characteristics offer promising directions for secure neural architectures. Future research directions include hardware implementation analysis through Intel’s Loihi neuromorphic processor, expanding privacy threat models, and integrating differential privacy mechanisms with evolutionary optimization in the TennLab framework. These investigations aim to further understand and enhance privacy preservation capabilities in neuromorphic architectures while maintaining their computational advantages.
Acknowledgements.
This work was funded in part by National Science Foundation through award CCF2319619 and in part by the CHIST-ERA grant TruBrain, by the UK’s Engineering and Physical Sciences Research Council (EPSRC) EP/Y03631X/1. The authors used AI-based tools, ChatGPT (OpenAI, 2024) and Claude (Anthropic, 2024) to revise the text in all Sections to correct any typos or grammatical errors.